PHYSICS 223 SP19 AA/AB QUIZ-3 Your Name Problem-1 Suppose that a muon is created in the upper atmosphere at an altitude of h-10.0km above the surface of the earth. Directly below on the surface of the earth is a scientist who observes the muon arrive and at her location and immediately decay (die) into an electron, electron-antineutrino and a muon-neutrino, μ + e + + , . The lifetime of a muon in its rest-frame is T.-2.20 x iO os. The rest-inass of a muon is 207 times the rest-mass of an electron, m-207me (a) What is the rest mass of the muon in units of MeV/c'? (b) What was the speed of the muon as it headed towards the scientist from the upper atmosphere?
PHYSICS 223 SP19 AA/AB Your Name: QUIZ-3 (c) What is the total energy of the muon in its own rest-frame? (d) What is the total energy of the muon in the scientist's rest-frame? madeptinta า)qqu 9dī moil 11itteibe alt dni ro, Inbaad 1i as n oum setr to te)qe 9tir 8nu 1'..I'// (d) (c) What is the kinetic energy of the muon in its own rest-frame? (f) What is the kinetic energy of the muon in the scientist's rest frame? Newtonian and relativistic kinetic energies of the Problem-2 What is the percent difference between the muon?
Answer #1

a)

Rest mass of an electron m_e=9.109390 \times 10^{-31}\ kg = 0.511 \ \frac{MeV}{c^2}

since the rest mass of a muon is m_{\mu}=207\ m_e we get

m_{\mu}=207 \times 0.511 = 105.77 \ \frac{MeV}{c^2}